OpenStudy (anonymous):

one more... Jason is a salesperson in the electronics department of a store. He is 40% short of his goal for the week. His goal is to earn $600. For each CD player he sells, he earns $48. How many CD players does he need to sell in order to meet his goal?

OpenStudy (anonymous):

600*0.4=240 240/48=CD now you can solve

OpenStudy (amistre64):

600/48 = 300/24 = 150/12 = 75/6 ; 12 1/2' 13 cds to get to 600 13 48 ---- 104 52 ---- 624 dollaars

OpenStudy (amistre64):

13(.4) = 5.2; he needs to sell 6 more cds..

OpenStudy (amistre64):

or; 600(.4) = 240/48= 120/24 = 60/12 = 5 cds

OpenStudy (amistre64):

240 = 5; 7*48 = 336 but; 7+5 cds only get him to $576